OpenStudy (anonymous):

A bicycle wheel is 9 inches in radius. The bicycle is travelling at miles per hour, what is the angular velocity, in revolutions per minute, of the wheel?

OpenStudy (tkhunny):

You'll need an actual number in there for the speed.

OpenStudy (anonymous):

Oops I left out the 25mph.

OpenStudy (tkhunny):

What's the circumference?

OpenStudy (anonymous):

Um.. 56.52?

OpenStudy (tkhunny):

\(2*\pi*9 = 56.55\) Not sure where you were rounding, but it was plenty short.

OpenStudy (anonymous):

Okay

OpenStudy (tkhunny):

Okay, now it's conversion time. \(\dfrac{25\;mile}{1\;hour}\cdot\dfrac{1\;Rev}{18\pi\;inch}\) What does that do for us?

OpenStudy (anonymous):

Converts it into revs.. What we need.

OpenStudy (tkhunny):

Good. Except we need RP(MINUTE). What else do we need in there?

OpenStudy (anonymous):

Minutes

OpenStudy (tkhunny):

Well, yes, but how shall it be introduced?

OpenStudy (anonymous):

Mm.. No clue.

OpenStudy (tkhunny):

How about \(\dfrac{1\;hour}{60\;min}\)?

OpenStudy (anonymous):

Oh yeah.. Then what would be the next step? 1min over..

OpenStudy (tkhunny):

You have hours in the denominator originally. In the new factor, hours need to be in the numerator to get rid of it. \(\dfrac{25\;mile}{1\;hour}\cdot\dfrac{1\;Rev}{18\pi\;inch}\cdot\dfrac{1\;hour}{60\;min}\) Is that something we can work with, or shall we do something about those mismatched linear measurements. Miles vs. inches? They are a little off.

OpenStudy (tkhunny):

25 miles = 25 * 5280 ft = 25 * 5280 * 12 in Are we looking familiar?